Another big couple of changes in the appearance of the experiment today. First, Andrei and Michele installed the radiating panels for the detector electronics' liquid cooling system (LCS) and connected all the plumbing. Then, in the early evening, a few of us stayed a little late and mounted the BTS, the large mylar-covered structure that surrounds the telescope and shields it from the sun. It's huge!
During the day, a bunch of other work was done as well. Matt and Franky fixed an issue with some of our DfMUX boards, Seth and Jeff fixed a CANBus issue (bad cable!), and Jeff additionally did some testing of his rebuilt HWP electronics crate (so far so good). A few people on the team left early to prepare for our first outdoor test tomorrow, a sensor calibration test.